OpenStudy (anonymous):

-86+2x=82-10x can someone show me the steps to complete this?

OpenStudy (anonymous):

\[-86+2x=82-10x\]Since 86 is negative we add it to both sides to move it..\[2x=168-10x\]Then, we have to add 10x to both sides..\[12x=168\]Finally, we divide both sides by 12 to get..\[x = 14\]

OpenStudy (anonymous):

-86 + 2x = 82 - 10x 2x + 10x = 82 + 86 12x = 168 x= 168/12 x= 14

OpenStudy (anonymous):

i wish i could give both of yall medals!